Marginal propensity to save is 0.3. Calculate MPC and the expenditure multiplier.
1
Expert's answer
2019-06-11T10:51:28-0400

Marginal propensity to save (MPS) is a part of the additional monetary unit received by the population, which increases the real cash income of the population and is aimed at additional savings.


The marginal propensity to save is nothing more than the inverse of the marginal propensity to consume (MPC), because in a simplified model of the economy, government activity and the influence of foreign trade are omitted, cash income will include savings and consumption, and therefore each monetary unit of income will disintegrate into consumption and savings.

MPS=1−MPC

MPC=1-MPS

MPC=1 - 0.3 = 0.7

The term “multiplier” means a multiplier, i.e. a number must be multiplied in order to get an increase in income. In turn, income is divided into savings and consumption.

MULT = 1/MPS or MULT = 1 / (1-MPS)

MULT = 1/0.3 = 3.33